# Dynamic Scheduling for Vehicle-to-Vehicle Communications Enhanced Federated Learning

### v1

@@ V-C Cooperative Transmission Problem
$V\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}\frac{1}{2}\kappa\beta\log_{2}\bigg{(}1+\frac{p^{\dagger}_{m}(t)|h_{m,r}(t)|^{2}}{\beta N_{0}}$
$+\sum_{n\in\mathcal{U}_{k}}\frac{u^{\dagger}_{n}(t)p^{\dagger}_{n}(t)|h_{n,r}(t)|^{2}}{\beta N_{0}}\bigg{)}-\frac{1}{2}\kappa q_{m}^{\text{SOV}}(t)p^{\dagger}_{m}(t)$
@@
$=V\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}\frac{1}{2}\kappa\beta\log_{2}\bigg{(}1+\frac{p^{\prime}_{m}(t)|h_{m,r}(t)|^{2}}{\beta N_{0}}$
$+\sum_{n\in\mathcal{U}_{k}}\frac{u^{\prime}_{n}(t)p^{\prime}_{n}(t)|h_{n,r}(t)|^{2}}{\beta N_{0}}\bigg{)}-\frac{1}{2}\kappa q_{m}^{\text{SOV}}(t)p^{\prime}_{m}(t)$
@@
     Update the virtual queue $\boldsymbol{q}(t)$ according to (19) and (20);
@@
When $\boldsymbol{u}(t)$ is given, the constraint (27c) becomes
$1+\frac{p_{m}(t)|h_{m,r}(t)|^{2}}{\beta N_{0}}+\sum_{n\in\mathcal{R}(t)}\frac{p_{n}(t)|h_{n,r}(t)|^{2}}{\beta N_{0}}$ (28)
$\leq 1+\frac{p_{m}(t)|h_{m,n}(t)|^{2}}{\beta N_{0}},\quad\forall n\in\mathcal{R}(t),$
$P4:$ $\underset{\boldsymbol{p}(t)}{\max}\ V\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}\frac{1}{2}\kappa\beta\log_{2}\bigg{(}1+\frac{p_{m}(t)|h_{m,r}(t)|^{2}}{\beta N_{0}}$
$+\sum_{n\in\mathcal{R}(t)}\frac{p_{n}(t)|h_{n,r}(t)|^{2}}{\beta N_{0}}\bigg{)}-\frac{1}{2}\kappa q_{m}^{\text{SOV}}(t)p_{m}(t)$
$-\sum_{n\in\mathcal{R}(t)}\frac{1}{2}\kappa q_{n}^{\text{OPV}}(t)p_{n}(t)$ (29)
[…]
V-D The Complete Algorithm
@@
Figure 3: The SUMO road network.
@@ VI Experiments
In this section, we evaluate the performance of the proposed VEDS algorithm. Firstly, it is compared with the benchmarks under different vehicle speeds $v$, approximation parameters $\alpha$, and weight $V$. Then, the proposed VEDS algorithm is evaluated for the CIFAR-10 image classification task [46]. Finally, the VEDS algorithm is applied to a real-world trajectory prediction dataset Argoverse [47] to showcase the value in practical vehicular applications.
@@
Energy Consumption Coefficient $10^{-28}$
Energy Constraints Random selected from $0.05$J to $0.1$J
VI-A Simulation setups
A road network is built based on SUMO [48], as shown in Fig. 3. An RSU is placed at the center of the road network. The vehicles move according to the Manhattan mobility model with a maximum speed of $v$ m/s, where $v$ is a variable for the experiments. For wireless communications, we adopt the V2X channel models in 3GPP TR 37.885 [11]. In the urban environment, the pathloss of the LOS and the NLOSv channels are given by $PL_{\text{LOS}}=38.77+16.7\log_{10}{d}+18.2\log_{10}{\gamma},$ where $d$ is the distance between two devices, $\gamma$ is the carrier frequency. The pathloss of the NLOS channel is specified by $PL_{\text{NLOS}}=36.85+30\log_{10}{d}+18.9\log_{10}{\gamma}.$ The simulation parameters are summarized in Table I.
Figure 6: The graph for the sigmoid function.
Figure 7: The graph for the derivative of the sigmoid function.
Figure 9: Total energy consumption of all vehicles per round under different weight parameter $V$.
For comparison, the following benchmarks are considered:
VI-A1 Optimal Benchmark
VI-A2 Dynamic algorithm with V2I-only communications (V2I-only)
VI-A3 Mobility and channel dynamic-aware FL (MADCA-FL)
This is a state-of-the-art VFL framework that considers the rapidly changing channel and vehicle mobility [7].
VI-A4 Static resource allocation and device scheduling algorithm (SA)
VI-B Performance of VEDS under different parameters
VI-B1 Impact of vehicle speed $v$
Firstly, we validate the performance of our algorithm under different vehicle speeds. We use the objective function of $P1$, i.e., the number of successful aggregations, as the performance metrics. As illustrated in Fig. 9, the number of successful aggregations of our framework initially increases and then decreases as the vehicle speed is adjusted from 0 (a stationary scenario) to $25$m/s, achieving $81.03\%$ of the optimal benchmark performance when $v=5$m/s. This performance increase at low speeds can be attributed to the mobility of vehicles allowing OPVs to enter the coverage of the RSU, while the SOVs largely remain within the RSU coverage area. If the vehicles move at high speed, the departure of some SOVs from RSU coverage results in deteriorated channel conditions. However, with the assistance of OPVs, these SOVs can still transmit the model back to the RSU. In comparison, the V2I-only framework and the MADCA-FL also consider vehicle mobility and exhibit certain robustness to changes in mobility. The SA framework, which employs static device scheduling, shows a significant performance decline in high-speed scenarios.
VI-B2 Impact of $\alpha$
We evaluate our proposed algorithm for different values of $\alpha$, as shown in Fig. 9. It is illustrated that as $\alpha$ increases from $10^{-2}$ to $10^{2}$, the number of successful aggregations first increases and then decreases, reaching a maximum when $\alpha$ is approximately equal to $2$. This is because as Theorem 2 suggests, when the parameter $\alpha$ is too small, the sigmoid function becomes overly smooth, leading to a suboptimal approximation of the indicator function (as shown in Fig. 9). On the other hand, when $\alpha$ is too large, the term $\psi(\alpha)$ diminishes, resulting in a loose bound in (22). Both scenarios adversely affect the overall performance of the algorithm. We also explain this phenomenon from a more intuitive perspective. As illustrated in Fig. 9, when $\alpha$ is small, the weight $\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}$ increases slowly with respect to $\zeta_{m}(t)$, the amount of transmitted model parameters. In this case, the algorithm tends to schedule vehicles evenly to balance their energy consumption. Consequently, it is possible that many vehicles have transmitted most of their model parameters but have not completed the upload. In the FL context, such a scenario is considered a transmission failure. When $\alpha$ is large, $\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}$ also increases slowly when $\zeta_{m}(t)$ is small, and thus, the aforementioned phenomenon persists, leading to suboptimal performance.
@@ VI-B3 Impact of $V$
Then, we evaluate our proposed VEDS algorithm for different weight parameters $V$. The number of successful aggregations and the energy consumption of all vehicles are shown in Fig. 9 and Fig. 9, respectively. It is illustrated that as $V$ increases from $10^{-2}$ to $10^{2}$, vehicles tend to consume more energy, which results in higher energy usage and a greater number of successful aggregations. When $V$ exceeds a threshold (around $V=1$), most vehicles use their maximum transmission power to upload their model, and the energy constraints are violated. Therefore, in practical systems, it is crucial to carefully choose the value of $V$ to ensure optimal training performance under energy constraints.
VI-C Evaluation on the CIFAR-10 dataset
Then, we evaluate the proposed VEDS algorithm on the CIFAR-10 dataset[46], which comprises $50000$ training images and $10000$ test images across ten categories. We consider both the independent and identically distributed (i.i.d.) and non-independent and identically distributed (non-i.i.d) settings. For the i.i.d. setting, the dataset is evenly divided into $40$ subsets, each containing samples from all 10 categories. For the non-i.i.d. setting, data samples are organized by category, and each vehicle holds a disjoint subset of data with samples from $2$ categories. Using the dataset, we train a convolutional neural network (CNN) with six convolutional layers. The learning rate is $0.1$, and the batch size is set to $32$.
The test accuracy of the VEDS algorithm compared with the benchmarks is illustrated in Fig. 12 (i.i.d.) and Fig. 12 (non-i.i.d.), where $\alpha=2,V=0.2$, and vehicle speed $v=10$ m/s. In the i.i.d. scenario, both VEDS and the benchmarks achieve high test accuracy. The VFL convergence speed of the VEDS algorithm closely approaches that of the optimal benchmark and is significantly higher than other benchmarks. Under the non-i.i.d. scenario, the convergence speed and the highest test accuracy of the VEDS algorithm are close to the optimal benchmark and exceed the other three benchmarks. After 1000 seconds of training, VEDS achieves a test accuracy of $62.45\%$, exceeding V2I-only, MADCA-FL and SA over $13.86\%$, $18.28\%$ and $41.50\%$. After 10000 seconds of training, the highest achievable accuracies are $79.41\%$ for the optimal benchmark, $78.80\%$ for VEDS, $76.33\%$ for V2I-only, $76.37\%$ for MADCA-FL, and $73.43\%$ for SA.

### v2

@@ V-C Cooperative Transmission Problem
$V\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}\frac{1}{2}\kappa\beta\log_{2}\bigg{(}1+\frac{p^{\dagger}_{m}(t)h_{m,r}(t)}{\beta N_{0}}$
$+\sum_{n\in\mathcal{U}_{k}}\frac{u^{\dagger}_{n}(t)p^{\dagger}_{n}(t)h_{n,r}(t)}{\beta N_{0}}\bigg{)}-\frac{1}{2}\kappa q_{m}^{\text{SOV}}(t)p^{\dagger}_{m}(t)$
@@
$=V\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}\frac{1}{2}\kappa\beta\log_{2}\bigg{(}1+\frac{p^{\prime}_{m}(t)h_{m,r}(t)}{\beta N_{0}}$
$+\sum_{n\in\mathcal{U}_{k}}\frac{u^{\prime}_{n}(t)p^{\prime}_{n}(t)h_{n,r}(t)}{\beta N_{0}}\bigg{)}-\frac{1}{2}\kappa q_{m}^{\text{SOV}}(t)p^{\prime}_{m}(t)$
@@
When $\boldsymbol{u}(t)$ is given, the constraint (28c) becomes
$1+\frac{p_{m}(t)h_{m,r}(t)}{\beta N_{0}}+\sum_{n\in\mathcal{R}(t)}\frac{p_{n}(t)h_{n,r}(t)}{\beta N_{0}}$ (29)
$\leq 1+\frac{p_{m}(t)h_{m,n}(t)}{\beta N_{0}},\quad\forall n\in\mathcal{R}(t),$
$P4:$ $\underset{\boldsymbol{p}(t)}{\max}\ V\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}\frac{1}{2}\kappa\beta\log_{2}\bigg{(}1+\frac{p_{m}(t)h_{m,r}(t)}{\beta N_{0}}$
$+\sum_{n\in\mathcal{R}(t)}\frac{p_{n}(t)h_{n,r}(t)}{\beta N_{0}}\bigg{)}-\frac{1}{2}\kappa q_{m}^{\text{SOV}}(t)p_{m}(t)$
$-\sum_{n\in\mathcal{R}(t)}\frac{1}{2}\kappa q_{n}^{\text{OPV}}(t)p_{n}(t)$ (30)
[…]
@@
     Update the virtual queue $\boldsymbol{q}^{\text{SOV}}(t)$ and $\boldsymbol{q}^{\text{OPV}}(t)$ according to (20) and (21);
@@
V-D The Complete Algorithm and Discussions
@@
To further elaborate on the advantages of VEDS in enhancing VFL performance, we discuss its impact from three key aspects. Firstly, dynamic scheduling enables real-time adaptation to vehicular mobility by scheduling OPVs with favorable channel conditions for relaying. By reducing the transmission distance between the SOV and the RSU, this strategy increases the transmission rate and the success rate of model uploads, especially in high-mobility environments. Secondly, energy balancing is achieved by cooperative model transmissions, reducing the power consumption required for direct SOV-to-RSU transmissions. Since received power diminishes rapidly with distance, utilizing OPVs for shorter-range relaying conserves energy, and prevents excessive battery depletion in SOVs. Thirdly, the use of DSTC in cooperative transmissions further enhances upload reliability by leveraging spatial diversity, allowing multiple OPVs to form a virtual multi-antenna system, and enhancing transmission reliability by mitigating channel fading. These mechanisms collectively improve efficiency, reliability, and sustainability in VFL.
@@ V-E Complexity Analysis
V-F Implementation Issues
To enable the efficient implementation of the VEDS algorithm in a real system, we provide practical considerations as follows.
V-F1 Synchronization in VFL Training
Maintaining synchronization in VFL systems is critical to ensure timely and efficient aggregation of local models from participating vehicles. Time synchronization protocols such as the network time protocol (NTP) can be introduced [50], ensuring all devices share a common time reference. NTP works by exchanging time-stamped messages between vehicles and the RSU, measuring the time delay during the transmission and adjusting the local clock of vehicles accordingly.
V-F2 Integration with Existing Vehicular Communication Protocols
To be implemented in real-world scenarios, the VEDS algorithm needs to be integrated with existing vehicular communication standards such as 5G NR-V2X [11] and LTE-V2X [10]. These protocols support both V2I and V2V communications, as well as the exchange of control information, which ensures the algorithm operates effectively within the constraints of current vehicular communication frameworks.
V-F3 V2V Communication Overhead
V2V communications introduce additional overhead, especially in scenarios with high vehicle mobility. Firstly, in dense vehicular networks, network congestion exacerbates communication delays. The VEDS algorithm mitigates this by prioritizing vehicle scheduling according to the amount of model parameters transmitted and channel conditions, to prevent simultaneous transmissions from multiple SOVs within the same area. Secondly, frequent link disconnections occur as vehicles continuously move in and out of the transmission range of each other. Fast connection mechanisms like proactive neighbor discovery [51] can be employed to enable vehicles to quickly detect and connect to new neighbors, reducing signaling overhead during reconnections. Finally, when implemented into practical systems, there might be potential interference from other vehicles not participating in the VFL process (e.g., those transmitting other signals). Our algorithm can be integrated with the sidelink semi-persistent scheduling (SPS) mechanism [52, 53]. The SPS mechanism in C-V2X randomly allocates one of the best resources based on channel sensing in the previous period, and keeps transmitting on the selected resource for certain times to minimize the interference and possibility of collisions. Our algorithm can schedule vehicles based on the resources allocated by SPS.
V-G Extension to Frequency Selective Channels
The VEDS algorithm introduced in the previous section assumes frequency non-selective channels. In this section, we extend its deployment to a frequency selective environment, where the additional dimension of channel variation across frequency should be considered. A feasible approach to extending the current design is to utilize the average channel gain for each vehicle, as presented in [43, 54, 55]. For an arbitrary channel $h_{m,n}(t)$, we divide it into $|\mathcal{N}_{m,n}|$ non-selective sub-channels, where $\mathcal{N}_{m,n}$ represents the set of sub-channel indices and $h_{m,n}^{(j)}(t)$ is the channel gain of the $j$-th sub-channel. The average channel gain is then expressed as:
$\bar{h}_{m,n}(t)=\frac{\sqrt{\sum_{j\in\mathcal{N}_{m,n}}\left(h_{m,n}^{(j)}(t)\right)^{2}}}{|\mathcal{N}_{m,n}|}.$
By substituting this average channel gain $\bar{h}_{m,n}(t)$ for $h_{m,n}(t)$, we can apply the VEDS algorithms in a frequency selective context.
@@ VI Experiments
In this section, we evaluate the performance of the proposed VEDS algorithm. Firstly, it is compared with the benchmarks under different system parameters. Then, the proposed VEDS algorithm is evaluated for the CIFAR-10 image classification task [56]. Finally, the VEDS algorithm is applied to a real-world trajectory prediction dataset Argoverse [57] to showcase the value in practical vehicular applications.
VI-A Simulation Setups
A road network is built based on SUMO [60], as shown in Fig. 3. An RSU is placed at the center of the road network. The vehicles move according to the Manhattan mobility model. At each intersection, a vehicle has a $50\%$ probability of going straight, and a $25\%$ probability of turning left or right. The vehicle speed is modeled by an intelligent driver model, with a maximum speed of $v^{\text{max}}$ m/s, where $v^{\text{max}}$ is a variable for the experiments. The speed of vehicle $m$ is controlled by the following equation:
$a_{m}=a^{\text{max}}\left[1-\left(\frac{v_{m}}{v^{\text{max}}}\right)^{4}-\left(\frac{d_{m}^{\prime}}{d_{m}^{*}}\right)^{2}\right],$
where $v_{m}$ is the instantaneous speed, $v^{\text{max}}$ is the maximum speed, $a_{m}$ is the instantaneous acceleration, $a^{\text{max}}$ is the maximum acceleration, $d_{m}^{\prime}$ is the distance between vehicle $m$ and the vehicle in front of it. $d_{m}^{*}$ is the desired dynamical distance, calculated as
$d_{m}^{*}=d^{\text{safe}}+v_{m}t^{\text{dst}}+\frac{v_{m}\cdot\Delta v_{m}}{2\sqrt{a^{\text{max}}a^{\text{brak}}}},$
where $d^{\text{safe}}$ is the safety distance between vehicle $m$ and the vehicle in front of it, $t^{\text{dst}}$ is the desired time headway that gives the minimum possible time to the vehicle directly in front of $m$, $\Delta v_{m}$ is the speed difference between $m$ and the vehicle in front of it, and $a^{\text{brak}}$ is a constant that defines the comfortable braking deceleration.
For wireless communications, we adopt the V2X channel models in 3GPP TR 37.885 [11]. In the urban environment, the pathloss of the LOS and the NLOSv channels are given by $PL_{\text{LOS}}=38.77+16.7\log_{10}{d}+18.2\log_{10}{\gamma},$ where $d$ is the distance between two devices, $\gamma$ is the carrier frequency. The pathloss of the NLOS channel is specified by $PL_{\text{NLOS}}=36.85+30\log_{10}{d}+18.9\log_{10}{\gamma}.$ The simulation parameters are summarized in Table I.
Figure 3: The SUMO road network.
For comparison, the following benchmarks are considered:
VI-A1 Optimal Benchmark
VI-A2 Dynamic algorithm with V2I-only communications (V2I-only)
Figure 6: The graph for the derivative of the sigmoid function.
Figure 8: Total energy consumption of all vehicles per round under different weight parameters $V$.
Figure 9: The average energy consumption of SOVs per round under different weight parameters $V$.
VI-A3 Vehicular edge FL (VEFL)
This is a state-of-the-art VFL framework that considers rapidly changing channels and vehicle mobility [41]. This framework considers that the vehicular network operates in a discrete time-slotted manner, and dynamically adjusts the resource allocation decisions in every slot.
VI-A4 Mobility and channel dynamic-aware FL (MADCA-FL)
This is another state-of-the-art VFL framework considering rapidly changing channels and vehicle mobility [7]. This framework makes scheduling and resource allocation decisions at the start of each round instead of in every slot.
VI-A5 Static resource allocation and device scheduling algorithm (SA)
VI-B Performance of VEDS under Different Parameters
VI-B1 Impact of maximum vehicle speed $v^{\text{max}}$
Firstly, we validate the performance of our algorithm under different maximum vehicle speeds. We use the objective function of $P1$, i.e., the number of successful aggregations, as the performance metrics. As illustrated in Fig. 9, the number of successful aggregations of our framework initially increases and then decreases as $v^{\text{max}}$ is adjusted from 0 (a stationary scenario) to $25$m/s, achieving $81.03\%$ of the optimal benchmark performance when $v^{\text{max}}=5$m/s. This performance increase at low speeds can be attributed to the mobility of vehicles allowing OPVs to enter the coverage of the RSU, while the SOVs largely remain within the RSU coverage area. If the vehicles move at high speed, the departure of some SOVs from RSU coverage results in deteriorated channel conditions. However, with the assistance of OPVs, these SOVs can still transmit the model back to the RSU. The comparison of VEDS with the V2I-only framework further highlights the advantages of V2V communications. At low speeds, where most SOVs remain within RSU coverage, VEDS achieves higher performance than V2I-only by dynamically scheduling OPVs with favorable channel conditions for relaying, reducing transmission distance and improving transmission rate. At high speeds, SOVs frequently leave RSU coverage before completing uploads, causing a significant performance drop in V2I-only. VEDS effectively utilizes OPVs to relay models, ensuring stable transmissions despite mobility. Among the other benchmarks, VEFL and the MADCA-FL frameworks also consider vehicle mobility and exhibit certain robustness to changes in mobility. The SA framework, which employs static device scheduling, shows a significant performance decline in high-speed scenarios.
VI-B2 Impact of $\alpha$
We evaluate our proposed algorithm for different values of $\alpha$, as shown in Fig. 9. It is illustrated that as $\alpha$ increases from $10^{-2}$ to $10^{2}$, the number of successful aggregations first increases and then decreases, reaching a maximum when $\alpha$ is approximately equal to $2$. This is because as Theorem 3 suggests, when the parameter $\alpha$ is too small, the sigmoid function becomes overly smooth, leading to a suboptimal approximation of the indicator function. On the other hand, when $\alpha$ is too large, the term $\psi(\alpha)$ diminishes, resulting in a loose bound in (23). Both scenarios adversely affect the overall performance of the algorithm. We also explain this phenomenon from a more intuitive perspective. As illustrated in Fig. 9, when $\alpha$ is small, the weight $\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}$ increases slowly with respect to $\zeta_{m}(t)$, the amount of transmitted model parameters. In this case, the algorithm tends to schedule vehicles evenly to balance their energy consumption. Consequently, it is possible that many vehicles have transmitted most of their model parameters but have not completed the upload. In the FL context, such a scenario is considered a transmission failure. When $\alpha$ is large, $\frac{d\sigma(\zeta_{m}(t))}{d\zeta_{m}(t)}$ also increases slowly when $\zeta_{m}(t)$ is small, and thus, the aforementioned phenomenon persists, leading to suboptimal performance.
@@ VI-B2 Impact of $\alpha$
Energy consumption coefficient $10^{-28}$ [58, 59]
Energy constraints per vehicles Selected from $0.02$J to $0.1$J
Total energy constraints per round $0.9$J
Length of time slot $0.1$s
@@ VI-B3 Impact of $V$
Then, we evaluate our proposed VEDS algorithm for different weight parameters $V$. The number of successful aggregations and the energy consumption of all vehicles are shown in Fig. 9 and Fig. 9, respectively. It is illustrated that as $V$ increases from $10^{-2}$ to $10^{2}$, vehicles tend to consume more energy, resulting in higher energy usage and more successful aggregations. When $V$ exceeds a threshold (around $V=1$), most vehicles use their maximum transmission power to upload their model, and the energy constraints are violated. Therefore, in practical systems, it is crucial to carefully choose the value of $V$ to ensure optimal training performance under energy constraints. Fig. 9 illustrates the average energy consumption of SOVs per round under different weight parameters $V$. Compared to V2I-only, while the total energy consumption is similar, VEDS achieves better energy balance by distributing transmission tasks among OPVs, reducing the average energy consumption per SOV.
VI-B4 Impact of Number of SOVs
Then, we evaluate the scalability of the VEDS algorithm as the number of SOVs in RSU coverage increases, as illustrated in Fig. 15. When the average number of SOVs in RSU coverage is below $12$, the number of successful aggregations of the proposed VEDS algorithm increases steadily as the number of SOVs grows, achieving close to optimal performance. When the average number of SOVs exceeds $12$, the increasing rate begins to slow down due the the limited communication and energy resources. Despite this, the proposed algorithm outperforms the benchmarks for all numbers of SOVs in the RSU coverage. The V2I-only, VEFL and MADCA-FL frameworks show moderate growth and exhibit performance gaps relative to VEDS. SA performs the worst due to its static scheduling mechanism, which fails to adapt to the increasing number of SOVs. These results demonstrate the scalability of the VEDS algorithm in accommodating a growing number of SOVs.
VI-B5 Impact of Model Size
We evaluate the performance of our algorithm under different model sizes, as shown in Fig. 15. For smaller model sizes (e.g., 1 or 2 million parameters), the number of successful aggregations of VEDS is close to the optimal value. As the model size increases, the number of successful aggregations gradually decreases due to the growing communication overhead. In contrast, other benchmarks show a sharper performance decline as the model size increases, particularly under larger model sizes, indicating less adaptability to increased model size.
Figure 10: Performance of the VEDS algorithm and the benchmarks under different numbers of SOVs.
Figure 11: Performance of the VEDS algorithm and the benchmarks under different model sizes.